# Entangled photon pair excitation and time-frequency filtered multidimensional photon correlation spectroscopy as a probe for dissipative exciton kinetics

### v1

@@ Appendix B Note on the observable: two-exciton population distribution
The observable is defined using the projection operator onto the desired two-exciton state, averaged over the time-dependent density operator. Assuming the perturbative regime of exciton-driving field interaction to hold, the density operator can be expanded to the fourth-order in incoming field-exciton interaction Hamiltonian to obtain
$O_{ff}(t)$ $=(-i)^{4}\text{Re}\int^{\infty}_{-\infty}d\tau_{4}\int^{\tau_{4}}_{-\infty}d\tau_{3}\int^{\tau_{3}}_{-\infty}d\tau_{2}\int^{\tau_{4}}_{-\infty}d\tau_{1}\Big\langle P_{+}(t)H_{\text{int},R}(\tau_{3})H_{\text{int},R}(\tau_{2})H_{\text{int},L}(\tau_{4})H_{\text{int},L}(\tau_{1})\Big\rangle$
@@
In the expression using superoperator notation [37, 12], the variables $\tau_{j}$ denoted the exciton-photon interaction times in the first line. It is converted into time-delays $t_{j}=\tau_{ij}=\tau_{i}-\tau_{j}$ in the subsequent line. The field expression correlate the exciton evolution via the time integrals. Expanding the exciton correlation function in the basis of one and two-exciton states, and carrying out the implicit averaging over the phonons lead to an computable expression of the signal which contains the exciton Green’s functions (defined in Eq. II). We further recast the expression in the frequency domain to obtain
$O_{ff}(t)$ $=2\,\text{Re}\Big\{\int\frac{d\omega_{4}}{2\pi}\int\frac{d\omega_{3}}{2\pi}\int\frac{d\omega_{2}}{2\pi}\int\frac{d\omega_{1}}{2\pi}e^{+i\omega_{3}\tau_{3}+i\omega_{4}\tau_{4}-i\omega_{2}\tau_{2}-i\omega_{1}\tau_{1}}\langle E^{\dagger}(\omega_{4})E^{\dagger}(\omega_{3})E(\omega_{2})E(\omega_{1})\rangle$
@@
In expression above is significantly more detailed. Here five types of phonon-averaged exciton Green’s functions, representing one, and two-exciton inter-manifold coherence, one and two-exciton transport and one-exciton intra-manifold coherence appear. They also bear accumulated frequencies of the driving field in their arguments. The first term represents the two-exciton coherence-mediated pathways, the second and fourth terms represents one-exciton transport-mediated pathways and the third and fifth terms represents the one-exciton coherence-mediated pathways. A more explicit form of the expression can be obtained if we plug in the expressions of the exciton Green’s functions. It gives
$O_{ff}(t)$ $\propto 2\mathrm{Re}\Big\{\int\frac{d\omega_{4}}{2\pi}\int\frac{d\omega_{3}}{2\pi}\int\frac{d\omega_{2}}{2\pi}\int\frac{d\omega_{1}}{2\pi}e^{+i\omega_{3}\tau_{3}+i\omega_{4}\tau_{4}-i\omega_{2}\tau_{2}-i\omega_{1}\tau_{1}}\langle E^{\dagger}(\omega_{4})E^{\dagger}(\omega_{3})E(\omega_{2})E(\omega_{1})\rangle$
@@
A compact version of this expression, using the abbreviation $I(\omega_{j};z_{ab})=(\omega_{j}+z_{ab})^{-1}$ is presented in Eqn. III.1. Proceeding via tedious yet straightforward frequency integrations yields the expression below
$O_{ff}(t)$ $\propto 2\text{Re}\Big[\exp{(z_{ff}t)}\Big(d_{fe^{\prime}}d_{e^{\prime}g}d_{fe}d_{eg}^{*}\times\big\langle E^{\dagger}(z_{fe^{\prime}}-z_{ff})E^{\dagger}(z_{fg}-z_{fe^{\prime}})E(-z_{eg}+z_{fg})E(z_{eg})\big\rangle$
@@
$+d_{fe^{\prime\prime}}d_{fe^{\prime\prime}}^{*}d_{e^{\prime}g}d_{eg}^{*}\times\big\langle E^{\dagger}(-z_{ef}+z_{ff})E(+z_{ee^{\prime}}-z_{ef})E^{\dagger}(z_{eg}-z_{ee^{\prime}})E(z_{eg})\big\rangle\Big)\Big]$ (27)
which is amenable to numerical simulation.
Appendix C Entangled photon correlation function
Here we present the numerical simulation of the square of the joint spectral amplitude function, for typical parameter regime used in the main simulation.
@@ Appendix C Entangled photon correlation function
Appendix D Temporal spectral filtering functions
We use the following temporal ${F}_{\bar{t}}(t)$, and spectral $F_{\bar{\omega}}(\omega)$ filtering functions, expressed as
$F_{\bar{t}}(t^{\prime})$ $=\theta(t^{\prime}-\bar{t})\exp({-\sigma_{\bar{t}}(t^{\prime}-\bar{t})})$ (28)
$F_{\bar{\omega}}(\omega^{\prime})$ $=i(\omega^{\prime}-\bar{\omega}+i\sigma_{\omega})^{-1}$ (29)
These functions have four tunable parameters: the centering variables, $\bar{t}$, $\bar{\omega}$, and the width $\sigma_{\bar{t}}$, $\sigma_{\bar{\omega}}$. The presented simulations in Fig. 7 illustrates their role in signal classification.
Appendix E Note on the observable: photon coincidence signal
Here, we present a brief outline of derivation of the bare photon coincidence counting observable (i.e., without the filtering stage), following the original theory developed in [13, 11]. The expression is already presented in Eq. III, and appears as
@@ Appendix E Note on the observable: photon coincidence signal
Here, the averaging of the field correlation term is carried out over the detector modes, and the auxiliary photon modes, where as the averaging of the exciton correlation term is carried out over two-exciton state following transport. In the presented theory, the leading order contributions to the signal arise from expansion of the time dependent density operator to the fourth order in emitted field. Assuming photon emission into vacuum modes, and following the standard algebraic manipulations one obtains
@@
This expression represents the exciton correlation function weighted by the density of detector modes. The correlation function can be further expanded in terms of individual pathways Fig. 2, yielding the expressions in Eq. III.3.

### v2

@@ Appendix C Note on the observable: two-exciton population distribution
Here, we present an outline for the derivation of the signal presented in Section III.1. The observable is defined using the projection operator onto the desired two-exciton state, averaged over the time-dependent density operator. Assuming the perturbative regime of exciton-driving field interaction to hold, the density operator can be expanded to the fourth-order in incoming field-exciton interaction Hamiltonian to obtain
$\rho_{ff}(t)$ $=(-i)^{4}\text{Re}\int^{\infty}_{-\infty}d\tau_{4}\int^{\tau_{4}}_{-\infty}d\tau_{3}\int^{\tau_{3}}_{-\infty}d\tau_{2}\int^{\tau_{4}}_{-\infty}d\tau_{1}\Big\langle P_{+}(t)H_{\text{int},R}(\tau_{3})H_{\text{int},R}(\tau_{2})H_{\text{int},L}(\tau_{4})H_{\text{int},L}(\tau_{1})\Big\rangle$
@@
In the expression, the variables $\tau_{j}$ and $t_{j}=\tau_{ij}=\tau_{i}-\tau_{j}$ denote the exciton-photon interaction times and time-delays, respectively. The four-point correlation function of the driving field correlate the four-point correlation function of the exciton operators via the time integrals. Expanding the exciton correlation function in the basis of one and two-exciton states, carrying out the averaging over the phonons and inserting the frequency domain variables lead the expression of the signal
$\rho_{ff}(t)$ $=2\,\text{Re}\Big\{\int\frac{d\omega_{4}}{2\pi}\int\frac{d\omega_{3}}{2\pi}\int\frac{d\omega_{2}}{2\pi}\int\frac{d\omega_{1}}{2\pi}e^{+i\omega_{3}\tau_{3}+i\omega_{4}\tau_{4}-i\omega_{2}\tau_{2}-i\omega_{1}\tau_{1}}\langle E^{\dagger}(\omega_{4})E^{\dagger}(\omega_{3})E(\omega_{2})E(\omega_{1})\rangle$
@@
The expression contains three types of exciton Green’s functions that describes inter-manifold coherence, intra-manifold coherence and transport. The first term represents the two-exciton coherence-mediated pathways, the second and fourth terms represents one-exciton transport-mediated pathways and the third and fifth terms represents the one-exciton coherence-mediated pathways. We also note that the Green’s functions contains frequencies of the driving field in their arguments. The expression can be simplified by plugging in the expressions of the exciton Green’s functions. It yields
$\rho_{ff}(t)$ $\propto 2\mathrm{Re}\Big\{\int\frac{d\omega_{4}}{2\pi}\int\frac{d\omega_{3}}{2\pi}\int\frac{d\omega_{2}}{2\pi}\int\frac{d\omega_{1}}{2\pi}e^{+i\omega_{3}\tau_{3}+i\omega_{4}\tau_{4}-i\omega_{2}\tau_{2}-i\omega_{1}\tau_{1}}\langle E^{\dagger}(\omega_{4})E^{\dagger}(\omega_{3})E(\omega_{2})E(\omega_{1})\rangle$
@@
A compact version of this expression, using the abbreviation $I(\omega_{j};z_{ab})=(\omega_{j}+z_{ab})^{-1}$ is presented in Eq. III.1. In the final step, tedious yet straightforward frequency integrations yields the expression below
$\rho_{ff}(t)$ $\propto 2\text{Re}\Big[\exp{(z_{ff}t)}\Big\{d_{fe^{\prime}}d_{e^{\prime}g}d_{fe}d_{eg}^{*}\times\big\langle E^{\dagger}(z_{fe^{\prime}}-z_{ff})E^{\dagger}(z_{fg}-z_{fe^{\prime}})E(-z_{eg}+z_{fg})E(z_{eg})\big\rangle$
@@
$+d_{fe^{\prime\prime}}d_{fe^{\prime\prime}}^{*}d_{e^{\prime}g}d_{eg}^{*}\times\big\langle E^{\dagger}(-z_{ef}+z_{ff})E(+z_{ee^{\prime}}-z_{ef})E^{\dagger}(z_{eg}-z_{ee^{\prime}})E(z_{eg})\big\rangle\Big\}\Big]$ (27)
This expression is valid for any general correlated photonic source and can be used in numerical simulation.
Appendix D Entangled photon correlation function
Here we examine the time-frequency properties entangled photon pairs. We plot the square of the joint spectral amplitude function, for a parameter regime comparable to that used in the simulation in Section III.
@@ Appendix D Entangled photon correlation function
Appendix E Note on the observable: two-photon counting signal
Here, we present a brief derivation of the photon coincidence counting signal. We follow the original steps outlined in [17, 15]. The final expression is already presented in Eq. III, and appears as
The bare two-photon counting signal is expressed as
@@ Appendix E Note on the observable: two-photon counting signal
The field correlation term is averaged over the detector modes in vacuum. The exciton correlation function is averaged over the the two-exciton states that corresponds to the population terms of the density operator following the transport. We assumed that the leading order contributions to the photon counting signal arise from expansion of the density operator to the fourth order in emitted photon field. The emitted field modes are expanded into plane wave mode and following the standard algebraic manipulations we obtain
@@
It highlights the fact that the four-point exciton correlation function weighted by the density of detector modes evaluated at emitted frequencies. The exciton correlation function can be expanded further to yield the Liouville space pathways depicted in the diagram in Fig. 3. The expansion also yields the expressions in  III.3.